OpenStudy (anonymous):

find cos theta if tan theta =-sqrt 3 over 3 and falls in quad 2. Will award medal if answer is given

OpenStudy (psymon):

Well, tangent is sine over cosine. We have the following potential fraction combinations in the 2nd quadrant: The following are given in the order of 2pi/3, 3pi/4, and 5pi/6 \[\frac{ \frac{ \sqrt{3} }{ 2 } }{ \frac{ -1 }{ 2 } }\] \[\frac{ \frac{ \sqrt{2} }{ 2 } }{ -\frac{ \sqrt{2} }{ 2 } }\] \[\frac{ \frac{ 1 }{ 2 } }{ \frac{ -\sqrt{3} }{ 2 } }\]

OpenStudy (anonymous):

so the answer is?

OpenStudy (psymon):

Well, the 3pi/4 answer is out. Look at the division of 2pi/3 and 5pi/6 and see. Not that hard to check which one

OpenStudy (anonymous):

the second one?

OpenStudy (psymon):

5pi/6 :3 So the second of the two I just mentioned, ya.

OpenStudy (anonymous):

so negative one half?

OpenStudy (anonymous):

i ned th answer if you want the medal

OpenStudy (psymon):

Its not about the medal. But the answer was 5pi/6. Since cosine is the denominator, then it would be -sqrt3/2

OpenStudy (anonymous):

okay so how did you get that?

OpenStudy (anonymous):

what did you divide to get that?

OpenStudy (psymon):

Well, there are only 3 possibilities in the 2nd quadrant. Since tangent is sine divided by cosine, I set up fractions of sin/cos for each of the 3 main points in quadrant 2. From there I see that the 2nd one I put, 3pi/4, would just give me -1. I need a sqrt3 in my answer, so it makes sense that it could be the top or bottom one. In terms of doing the division, I would do this: \[\frac{ \frac{ 1 }{ 2 } }{ \frac{ -\sqrt{3} }{ 2 } } = \frac{ 1 }{ 2 } * \frac{ -2 }{ \sqrt{3} } = \frac{ -1 }{ \sqrt{3} } = \frac{ -\sqrt{3} }{ 3 }\]

OpenStudy (psymon):

At 5pi/6, the sine value is 1/2 and the cosine valueis -rt3/2. dividing those in that fashion eventually leads me to the value you're looking for.

OpenStudy (anonymous):

Thanks man

OpenStudy (psymon):

Np ^_^

OpenStudy (anonymous):

what about find cos theta if sin theta=-2/3 and falls in quadrant 4

OpenStudy (anonymous):

@Psymon

OpenStudy (psymon):

 I can show this as a triangle drawn in the 4th quadrant. Because sin represents the opposite side of a triangle over the hypotenuse, I can assign these two values to the appropriate sides of this triangle. I now need to solve for the missing side, which can be done through pythagorean theorem. Cosine is positive in the 4th quadrant, so there is no need to keep the negative sign with my answer, which is why I did not type it. So once you use the pythagorean theorem to find the missing side, use the fact that cosine represents the adjacent side of atriangle over the hypotenuse and this will give you the cosine valueyou need.

OpenStudy (anonymous):

okay so whats next?

OpenStudy (psymon):

Do you know how to solve for the missing side?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

can you show me really really quick bro?

OpenStudy (psymon):

Well, we need pythagorean theorem. The pythagorean theorem is a^2 + b^2 = c^2  so what I have is a missing in this case. If I want to solve for a, I can rearrange the pythagorean theorem and say that a^2 = c^2 - b^2. Therefore: \[a = \sqrt{(3)^{2}-(2)^{2}}\]

OpenStudy (anonymous):

so it'll be positive \[\frac{ \sqrt{5} }{ 3 }\] or \[\frac{ 2\sqrt{5} }{ 5 }\]

OpenStudy (anonymous):

i want to say the first one @Psymon

OpenStudy (psymon):

Yep, the first one would be correct : )

OpenStudy (anonymous):

okay so lets say same equation but it was in quadrant 3? would it only make it negavtive or would it be a different equation?

OpenStudy (anonymous):

@Psymon

OpenStudy (psymon):

It would simply be negative.

OpenStudy (anonymous):

ohhhh makes sense then so it would be a whole different answer?

OpenStudy (psymon):

Pretty much. But only because of the fact that cosine is negative in the 3rd quadrant.

OpenStudy (anonymous):

okay thank you soo much

OpenStudy (psymon):

Yep yep :3

OpenStudy (anonymous):

You're great help. I posted something about you sooo thanks and if I have anyhting I need help on. I'm coming to you bro.

OpenStudy (anonymous):

You really are the best @Psymon